OpenStudy (anonymous):

Find the lateral area of a conical grain storage tank that has a height of 30 meters, a diameter of 20 meters, and a slant height of 32 meters. Round the answer to the nearest square meter.

jhonyy9 (jhonyy9):



jhonyy9 (jhonyy9):

than you ,,open" a con on on the line of the slant height result a semicircle with radius equal the length of the slant height so this mean that in case of your exercise will result a semicircle with radius = 32 m area of a circle is equal pi*r^2 so than area of a semicircle will be area of circle divide by 2 so in case of your exercise will get area con = 3,14 *32 /2 = 3,14 * 16 = 50,24 so this mean that lateral area of this conical grain storage tank will be 50 square meter hope so much that is understandably sure right easy good luck bye